### v1

@@ Abstract
We show that there are $f$-vectors of balanced simplicial complexes giving a source of simplicial complexes exhibiting Boolean a decomposition similar to a geometric Lefschetz decomposition. The objects we are working with are $h$-vectors of flag spheres and balanced simplicial complexes whose $f$-vectors are equal to them. This builds on work of Nevo–Petersen–Tenner on a conjecture of Nevo–Petersen that the gamma vector of an odd-dimensional flag sphere is the $f$-vector of a balanced simplicial complex (which was for barycentric subdivisions by Nevo–Petersen–Tenner). We can connect our decomposition to positivity questions on reciprocal/palindromic polynomials associated to flag spheres and geometric questions motivating them. In addition, we note that that the degrees in the Lefschetz-like decomposition are not halved unlike the usual $h$-vector setting.
@@ Introduction
The objects which we will discuss are combinatorial objects with $f$-vectors that seem to have some kind of geometric structure. They come from flag spheres, which are simplicial spheres with special geometric properties. For example, they are both involved in questions about a special case of the Hopf conjecture for certain piecewise Euclidean manifolds in the sense of Gromov’s CAT(0) inequality [8] and postivity properties of reciprocal/palindromic polynomials lying between unimodality and real-rootedness observed in many combinatorial settings [3].
@@
A simplicial complex is flag if every minimal nonface has size 2. Equivalently, it is the clique complex of a graph.
@@ 1 Background
A simplicial complex is flag if every minimal nonface has size 2. Equivalently, it is the clique complex of a graph.
@@
    Let $\Delta$ be a $(d-1)$-dimensional Cohen–Macaulay simplicial complex. Setting $M=k[\Delta]$ and $\theta=(\theta_{1},\ldots,\theta_{d})$ to be a linear system of parameters, Part 1 implies that $k[\Delta]$ is a free module over $k[\theta]$ with basis $\eta_{1},\ldots,\eta_{t}$ if and only if $\eta_{1},\ldots,\eta_{s}$ is a $k$-basis for the Artinian reduction $A^{\cdot}(\Delta)\coloneq k[\Delta]/\theta k[\Delta]$ of the Stanley–Reisner ring $k[\Delta]$.
@@ 2.1 Boolean decompositions and $f$-vectors of balanced simplicial complexes
  • If we take $\Gamma$ to be the compression complex for $h(\Delta)$, its $(k-1)$-dimensional faces are the first $h_{k}(\Delta)$ elements of $\binom{\mathbb{N}}{k}$ (in reverse lexicographical order).
@@
Note that the partition of unity map applies to $20\leq k\leq d-$ 2. The generators of these parts consist of faces $H\in\Gamma$ of dimension $\leq d-3$. For these parts, we can make use of the $F_{e}\cup G_{e}$ representation of $H$ on the component corresponding to an edge $e\in\Delta$. Since the partition of unity map is still an injection after decreasing the degree to a smaller positive number, there is a unique $F\in\Gamma$ mapping to a given tuple $(F_{e})_{\begin{subarray}{c}e\in\Delta\\ |e|=2\end{subarray}}$ and a unique $G\in\Gamma$ mapping to a given tuple $(G_{e})_{\begin{subarray}{c}e\in\Delta\\ |e|=2\end{subarray}}$. We can set these $F$ and $G$ to be our initial and Boolean components respectively. If the map is simply trivial, it respects the inclusions. Since we use the compression complex for everything, the coordinates on for each edge $e\in\Delta$ are the same. Since that a subset of a face $H\in\Gamma_{e}$ is also a face of $\Gamma_{e}$, the $F_{e}$ and $G_{e}$ are nonzero at the components $e\in\Delta$ where $H$ is nonzero. The main thing to note is that $F_{e}$ and $G_{e}$ can be taken as is since the induced partition of unity map for the compression complexes is still the identity on components where it is nonzero and 0 on the other ones. Taking compression complexes, we can take the $(d-4)$-skeleton of each $\Gamma_{e}$ (i.e. everything except for the unique facet) to be a face of the $(d-2)$-skeleton of $\Gamma$ excluding the unique facet of $\Gamma$. In other words, we will take $\Gamma_{e}$ to be a collection of subcomplexes $\Gamma_{e}\leq\Gamma$ where each face of $\Gamma$ of dimension $\leq d-3$ is a face of $\Gamma_{e}$ for some $e\in\Delta$. For each such face of $\Gamma$, the injective form of the partition of unity map says that there is a unique tuple of such $\Gamma_{e}$ associated to each face $H\in\Gamma_{e}$.
@@
It remains to show that $\dim S\leq\frac{d}{2}-1$, where $S$ is the dimension of the subcomplex $S$ defined as the (vertex) restriction of $\Gamma$ to faces that are disjoint from the unique facet $[d]$ of $\Gamma$. Given a face of $\Gamma$, we can partition its vertices into a face $F\in S$ and $G\in 2^{[d]}$. For the unique facet $[d]$ of $\Gamma$, the former is the empty set and the latter is all of $[d]$. Now consider the ridges ($(d-2)$-dimensional faces) of $\Gamma$. We note that removing a single vertex yields a $(d-3)$-dimensional face of $\Gamma$. If the vertex $p$ is removed from the first part, then we have $F\setminus p\in S_{e}$ for some edge $e\in\Delta$ and $|F\setminus p|\leq\dim S_{e}+1\leq\frac{d}{2}-1\Longrightarrow|F|\leq\frac{d}{2}$. Suppose that the removed vertex $p$ is from the Boolean part. Then, we have that $F\in S_{e}$ for some edge $e\in\Delta$ and $|F|\leq\dim S_{e}+1\leq\frac{d}{2}-2+1=\frac{d}{2}-1$. As for smaller dimensional faces of $\Gamma$, they belong to $\Gamma_{e}$ for some edge $e\in\Delta$ and $F_{e}\in S_{e}$ gives the $S$-component since being disjoint from $[d]$ also means being disjoint from $[d-2]\subset[d]$.

### v2

@@ Abstract
We show that there are $f$-vectors of balanced simplicial complexes giving a source of simplicial complexes exhibiting a Boolean decomposition similar to a geometric Lefschetz decomposition. The objects we are working with are $h$-vectors of flag spheres and balanced simplicial complexes whose $f$-vectors are equal to them. This builds on work of Nevo–Petersen–Tenner on a conjecture of Nevo–Petersen that the gamma vector of an odd-dimensional flag sphere is the $f$-vector of a balanced simplicial complex (which was shown for barycentric subdivisions by Nevo–Petersen–Tenner). We can connect our decomposition to positivity questions on reciprocal/palindromic polynomials associated to flag spheres and geometric questions motivating them. In addition, we note that the degrees in the Lefschetz-like decomposition are not halved unlike the usual $h$-vector setting.
@@ Introduction
The objects which we will discuss are combinatorial objects with $f$-vectors that seem to have some kind of geometric structure. They come from flag spheres, which are simplicial spheres with special geometric properties. For example, they are both involved in questions about a special case of the Hopf conjecture for certain piecewise Euclidean manifolds in the sense of Gromov’s CAT(0) inequality ([8], Section 4.2.C on p. 122 of [17]) and postivity properties of reciprocal/palindromic polynomials lying between unimodality and real-rootedness observed in many combinatorial settings [3].
@@
A simplicial complex is flag if every minimal nonface has size 2. In other words, a subset $A\subset V(\Delta)$ of the vertex set $V(\Delta)$ is a face of $\Delta$ if and only if $e\in A$ for every 2-element subset $e\subset A$. Equivalently, it is the clique complex of a graph.
@@
Definition 0.5.
(Definition 4.1 on p. 95 of [31])
Let $\Delta$ be a $(d-1)$-dimensional simplicial complex on the vertex set $V$. We say that $\Delta$ is balanced if there is a map $\kappa:V\longrightarrow[d]$ such that if $\{x,y\}$ is an edge of $\Delta$, then $\kappa(x)\neq\kappa(y)$.
Equivalently, if we think of $\kappa$ as a coloring of $V$ with the colors $1,\ldots,d$, then every face of $\Delta$ has all its vertices colored differently.
@@ 1 Background
A simplicial complex $\Delta$ is flag if every minimal nonface has size 2. In other words, a subset $A\subset V(\Delta)$ of the vertex set $V(\Delta)$ is a face of $\Delta$ if and only if $e\in A$ for every 2-element subset $e\subset A$. Equivalently, it is the clique complex of a graph.
@@
Next, we mention constructions and properties of flag spheres (and flag simplicial complexes in general) which will be used for studying local-global properties and inductive arguments involving the dimension.
Definition 1.3.
(p. 60 of [31], Definition 5.3.4 on p. 232 of [5]) Given a simplicial complex $\Delta$, its link over a face $F\in\Delta$ is the subcomplex
$\operatorname{lk}_{\Delta}(F)\coloneq\{G\in\Delta:F\cup G\in\Delta,F\cap G=\emptyset\}.$
These are the faces of $\Delta$ formed by removing $F$ from faces of $\Delta$ containing $F$.
We first list a special local property of flag spheres before listing the codimension 2 subcomplexes which will be used for inductive arguments.
@@
Proposition 1.5.
Given a flag homology sphere $\Delta$, a link $\operatorname{lk}_{\Delta}(e)$ over an edge $e\in\Delta$ is also a flag homology sphere.
Flagness and being a homology sphere is preserved under taking links over vertices (Definition 1.2.1 on p. 271 of [15] and Theorem 5.1 on p. 65 of [31]).
@@
    Let $\Delta$ be a $(d-1)$-dimensional Cohen–Macaulay simplicial complex. Setting $M=k[\Delta]$ and $\theta=(\theta_{1},\ldots,\theta_{d})$ to be a linear system of parameters, Part 1 implies that $k[\Delta]$ is a free module over $k[\theta]$ with basis $\eta_{1},\ldots,\eta_{s}$ if and only if $\eta_{1},\ldots,\eta_{s}$ is a $k$-basis for the Artinian reduction $A^{\cdot}(\Delta)\coloneq k[\Delta]/\theta k[\Delta]$ of the Stanley–Reisner ring $k[\Delta]$.
@@ 2.1 Boolean decompositions and $f$-vectors of balanced simplicial complexes
  • If we take $\Gamma$ to be the compression complex for $h(\Delta)$, its $(k-1)$-dimensional faces are the first $h_{k}(\Delta)$ elements of $\binom{\mathbb{N}}{k}$ (in reverse lexicographical order). This can be used to $\Gamma_{e}$ interpret as a subcomplex of $\Gamma$. Note that other vertex orderings can also be used.
@@
  • The idea is to think about a cover of faces of $\Gamma$ by subcomplexes $\Gamma_{e}$ where $\Gamma$ and $\Gamma_{e}$ are analogous to $\Delta$ and $\operatorname{lk}_{\Delta}(e)$ in the partition of unity map. By a “cover”, we mean that every face of $\Gamma$ is a face of $\Gamma_{e}$ for some $e$. To be more precise, consider the surjective form of the partition of unity map and think about the basis elements of $A^{k}(\operatorname{lk}_{\Delta}(e))$ as faces of a subcomplex $\Gamma_{e}\leq\Gamma$ corresponding to subcomplexes forming a cover of $\Gamma$. Then, dualize the map to get the injective form.
@@
  • The same thing applies to a correspondence with pairs $(P,Q)$ of “independent” sets $P$ and $Q$ satisfying the properties above for $F$ and $G$ where $P$ and $Q$ are not necessarily the same kind of object.
@@
Note that the partition of unity map applies to $20\leq k\leq d-$ 2. The generators of these parts consist of faces $H\in\Gamma$ of dimension $\leq d-3$. For these parts, we can make use of the $F_{e}\cup G_{e}$ representation of $H$ on the component corresponding to an edge $e\in\Delta$. Since the partition of unity map is still an injection after decreasing the degree to a smaller positive number, there is a unique $F\in\Gamma$ mapping to a given tuple $(F_{e})_{\begin{subarray}{c}e\in\Delta\\ |e|=2\end{subarray}}$ and a unique $G\in\Gamma$ mapping to a given tuple $(G_{e})_{\begin{subarray}{c}e\in\Delta\\ |e|=2\end{subarray}}$. We can set these $F$ and $G$ to be our initial and Boolean components respectively. If the map is simply trivial, it respects the inclusions. If we use the compression complex for everything, the coordinates on for each edge $e\in\Delta$ are the same. In this case, the inclusion wouldn’t be a literal identity map but a correspondence. For example, different $(d-3)$-dimensional faces of $\Gamma$ would come from faces of different $\Gamma_{e}$. Alternatively, we can use a different ordering for the compression complex to get something close to a literal inclusion map instead of an injection.
As mentioned earlier, the point is that we can think about this as corresponding a cover of $\Gamma$ by subcomplexes $\Gamma_{e}$. Since a subset of a face $H\in\Gamma_{e}$ is also a face of $\Gamma_{e}$, the $F_{e}$ and $G_{e}$ are nonzero at the components $e\in\Delta$ where $H$ is nonzero. To take $G_{e}$ as part of the “Boolean” part after the embedding, we can use the compression complexes for $\Gamma_{e}$ using vertex orders (for reverse lexicographical order) where the first $d-2$ elements are given by different $(d-2)$-element vertex subsets of the unique facet $[d]\in\Gamma$ but repeated pairs of vertices in different vertices have the same order. This is similar to excluding colors not used by $\Gamma_{e}$ covering $\Gamma_{e}$ when we’re given a particular $d$-coloring of the vertices of $\Gamma$. The main thing to note is that $F_{e}$ and $G_{e}$ can be taken as is since the induced partition of unity map for the compression complexes is still the identity on components where it is nonzero and 0 on the other ones. Note that there are still $2^{d-2|F|-2}$ choices of $G$ for a given $F$ since the map is injective. We can also think about the pairing comment on $(P,Q)$ above. Taking compression complexes, we can take the $(d-4)$-skeleton of each $\Gamma_{e}$ (i.e. everything except for the unique facet) to be a face of the $(d-2)$-skeleton of $\Gamma$ excluding the unique facet of $\Gamma$. In other words, we will take $\Gamma_{e}$ to be a collection of subcomplexes $\Gamma_{e}\leq\Gamma$ where each face of $\Gamma$ of dimension $\leq d-3$ is a face of $\Gamma_{e}$ for some $e\in\Delta$. For each such face of $\Gamma$, the injective form of the partition of unity map says that there is a unique tuple of such $\Gamma_{e}$ associated to each face $H\in\Gamma_{e}$.
@@
It remains to show that $\dim S\leq\frac{d}{2}-1$, where $S$ is the dimension of the subcomplex $S$ defined as the (vertex) restriction of $\Gamma$ to faces that are disjoint from the unique facet $[d]$ of $\Gamma$. Given a face of $\Gamma$, we can partition its vertices into a face $F\in S$ and $G\in 2^{[d]}$. For the unique facet $[d]$ of $\Gamma$, the former is the empty set and the latter is all of $[d]$. Now consider the ridges ($(d-2)$-dimensional faces) of $\Gamma$. We note that removing a single vertex yields a $(d-3)$-dimensional face of $\Gamma$. If the vertex $p$ is removed from the first part, then we have $F\setminus p\in S_{e}$ for some edge $e\in\Delta$ and $|F\setminus p|\leq\dim S_{e}+1\leq\frac{d}{2}-1\Longrightarrow|F|\leq\frac{d}{2}$. Suppose that the removed vertex $p$ is from the Boolean part. Then, we have that $F\in S_{e}$ for some edge $e\in\Delta$ and $|F|\leq\dim S_{e}+1\leq\frac{d}{2}-2+1=\frac{d}{2}-1$.